Find three consecutive positive odd integers such that the product of the first two is eight more than five times the third.
Answer by ankor@dixie-net.com(22740)

Find three consecutive positive odd integers
x, (x+2), (x+4)
such that the product of the first two is eight more than five times the third.
x(x+2) = 5(x+4) + 8
x^2 + 2x = 5x + 20 + 8
combine like terms on the left to form a quadratic equation
x^2 + 2x - 5x - 28 = 0
x^2 - 3x - 28 = 0
Factor to
(x-7)(x+4) = 0
the positive solution is what they want here
x = 7, 9, 11 are the three integers
:
:
See if that works out in the statement:
"the product of the first two is eight more than five times the third."
7(9) = 5(11) + 8
63 = 55 + 8